### v1

@@ A.3 Reliability of IPCA
We tested the IPCA method on a fiducial dataset in order to illustrate its effectiveness at producing unbiased images of extended circumstellar signals, hence validate all our conclusions regarding the flux and morphology of circumstellar signals recovered in our IPCA images of PDS 70. For this test, we considered the only other JWST/NIRCam dataset obtained in the F187N filter with the NRCB1 detector and SUB64P subarray publicly available at the time of this study: an observation of reference star HD 135067 obtained on 2 February 2023 (Program 1902). This source does not have any known resolved circumstellar emission. We downloaded the calibrated i2d images directly from the MAST archive. While the number of integrations and the number of groups per integration were similar to the PDS 70 observation, the main difference was the adoption of a 3-point dither pattern strategy, instead of a 5-point dither pattern for the observation of PDS 70. This results in a larger sensitivity to spatial undersampling effects, expected to be particularly prominent near the core of the PSF. As a consequence, the results of the test presented in this section should be considered as a conservative lower limit on the actual expected performance of IPCA on the PDS 70 dataset.
@@
We designed a toy model for the circumstellar signals of PDS 70 composed of an outer disk, an inner disk, a spiral-like signal, and three point source injections corresponding to protoplanets and , and candidate . For the outer disk, we considered the optimal model found in Appendix B. For the inner disk, we considered similar parameters as presented in Keppler et al. (2018), with the peak of the emission concentrated within 5 au. The two disk components are shown in Fig. 6a. For the spiral-like signal, we considered a similar trace as observed in our F187N images of PDS 70 after subtraction of the optimal outer disk model (Fig. 6b). For the protoplanets, we considered similar positions and contrasts as inferred with NEGFC in Appendix C (Fig. 6c). The sum of all components of the model is shown in Fig. 6d. To create and inject the fiducial model into the reference cube, we relied on routines from the fm (forward modeling) subpackage of VIP, in particular its implementation of the Grenoble RAdiative TransfER tool (Augereau et al. 1999; Milli et al. 2019) for the inner disk model, its trace and fake companion injection routines. Before injection in the reference cube data, the model was scaled to a similar contrast with respect to the reference star as the circumstellar signals with respect to PDS 70. The signals were injected at two different angles separated by 5.0 deg in the two sets of 3 dithered images (i.e., the same roll angle difference as in the PDS 70 observation).
@@
Finally, we ran IPCA on this fiducial dataset using the same reduction parameters as used to produce the F187N images of PDS 70. The algorithm converged within 400 iterations, resulting in the image shown in Fig. 6e. To facilitate the comparison with the ground truth injected model, we highlight the location of the injected protoplanets with dashed circles. Apart from stellar residual artefacts, which can likely be assigned to small differences in the core of the PSF due to undersampling effects, we note a satisfactory recovery of both the morphology and flux level of the injected circumstellar signals. This casts confidence on the results obtained with IPCA on the PDS 70 dataset reported in this paper.
@@ Appendix B Optimal disk model found with NEGFD
To be able to extract unbiased photometry for the protoplanets, one needs to remove the expected contribution of the disk. Our goal in this work is not a full modeling of the disk, as this would involve a combined SED and disk image fit (e.g., Keppler et al. 2018; Portilla-Revelo et al. 2022, 2023). Therefore, we decided to rely on the latest radiative transfer models of the disk produced with MCMax3D (Min et al. 2009) and presented in Portilla-Revelo et al. (2022, 2023), as they match these combined constraints, and only allow these models to vary to a small extent – for them to still be compatible with ALMA and SPHERE polarized intensity constraints. Namely, we allow (i) the minimum grain size in the grain size distribution to vary between 0.001 and 0.05 m; (ii) the settling parameter to vary between 0.001 m and 0.01 m; (iii) spatial and flux scaling of the model image to vary within a small range around unity; and (iv) small linear (sub-pixel) and azimuthal (sub-degree) shifts with respect to the center of the image, as these parameters are essentially constrained by imaging. Small variations of these parameters can lead to noticeable changes in the model images, hence to significant residuals after subtraction of the model. We produced a grid of 10 disk models in (two explored values) and (five explored values), and searched for the optimal model with a Nelder-Mead algorithm, by interpolating model images in log-space and including free parameters for scaling and shifts. The model images were produced at the same pixel scale as the F187N and F480M images, and smoothed by convolution with the observed PSF. The optimal model was then found by minimizing absolute residuals after subtraction of the model to the observed images. This procedure, which we refer to as the Negative Fake Disk technique (NEGFD) is implemented in VIP, as of version 1.6.0.
@@
The optimal disk model was found by minimizing the sum of absolute intensity residuals in a binary mask which encompasses the location of the outer disk, while excluding the location of both planets (2-FWHM aperture exclusion masks). We considered two potential masks. The first one includes the whole forward scattering (i.e., brighter) side of the disk, and the second one only includes the southwest part of the disk, anticipating excess signals towards the west (near planet c) and northwest part of the disk (spiral-like feature) based on previous images of the disk (Wang et al. 2020; Juillard et al. 2022). As our tests using the first mask led to a mix of strong positive and negative residuals, we only consider the results obtained by minimizing residuals in the second mask, shown in Fig. 7c, in the rest of this work. In practice, we identified the optimal radiative transfer disk model using the F480M data because of the higher signal-to-noise ratio of the disk in these data. The optimal and values were found to be 0.001 m and 0.01, respectively, meaning that the optimal is in-between the radiative transfer models considered in Portilla-Revelo et al. (2022) and Portilla-Revelo et al. (2023). These parameters were then used to make a disk model prediction at 1.87 m. The best F187N and F480M models are shown in Fig. 7a and b, respectively.
@@
Figs. 8b and e show the roll subtraction images obtained after subtraction of the optimal disk model found by NEGFD from the original images (i.e., before PSF subtraction), for the F187N and F480M datasets, respectively. These are compared to the original roll subtraction images (Figs. 8a and d) using the same intensity scale. The disk subtraction performs well, in particular for the southwest part of the disk, where the residuals reach a similar level as the residual noise level in the image. Subtraction of the optimal disk model clearly highlights the presence in the F480M image of the arm-like feature characterized in Juillard et al. (2022) in VLT/SPHERE images. This feature is indicated with a thick arrow. We also checked how a similar NEGFD procedure performed on the IPCA images obtained with both filters, fixing the and to the optimal values found by NEGFC+roll subtraction. The results are shown in Fig. 1b and e, and discussed in Sec. 3.1.
@@ Appendix C NEGFC retrieval of planet parameters
The Negative Fake Companion (NEGFC) technique consists in finding the optimal parameters (radial separation, position angle and contrast with respect to the star) of a directly imaged companion through the injection of fake companions with a negative flux in the input image cube (i.e., before post-processing), such that residuals are minimized in the post-processed image around the location of the planet. This forward modeling approach is necessary to avoid the parameter estimates for the companion to be affected by self- and over-subtraction effects inherent to the post-processing algorithm used.
We used the Nelder-Mead minimization algorithm implemented in VIP, and described in Wertz et al. (2017) to estimate the parameters for the protoplanets. This NEGFC implementation offers different options in terms of figure of merit to be used in the processed image after injection of negative fake companions to identify the optimal radial separation , position angle PA and contrast of a given planet, namely minimizing (i) the sum of absolute intensity residuals in an aperture encompassing the companion, (ii) the standard deviation of intensity residuals in an aperture encompassing the companion, or (iii) the sum of absolute determinant values of the Hessian matrix calculated for each of the pixels surrounding the planet location. The latter option is equivalent to minimizing the local absolute curvature, and is more appropriate for the extraction of point source astrometry and photometry in the presence of underlying extended signals (e.g., Quanz et al. 2015). We implemented it for this work, and made it available in VIP as of version 1.6.0. Hereafter, we detail two different approaches involving NEGFC which we tested and compared to retrieve optimal astrometry and photometry for the protoplanets.
@@ C.1 Classical forward-modeling NEGFC
We first considered the classical forward-modeling NEGFC approach, as described in Lagrange et al. (2010) or Wertz et al. (2017), combined with roll subtraction, hereafter referred to as NEGFC+roll. We applied NEGFC+roll to the image cube where the optimal disk model found by NEGFD was removed (Appendix B). Regarding the choice of figure of merit, we note that different figures of merit lead to more or less reliable retrieved planet parameters depending on the SNR of the companion, the nature of the local noise, and contamination by residual extended signals (e.g., inner disk or spiral accretion stream signal). We therefore determined the optimal figure of merit on a case by case basis for each planet in each of our two datasets, as follows. We used NEGFC to derive first estimates on the parameters of the companion with each of the three figures of merit, considering two subcases for the sum and standard deviation, corresponding to their calculation in either 1-FWHM or 2-FWHM size apertures, and three subcases for the Hessian-matrix determinant figure of merit, corresponding to its calculation with set to 1, 2 or 3. Then we injected 360 fake (positive) companions with the radial separation and contrast inferred by NEGFC, 1 apart from each other, in separate copies of the original datasets, and retrieved individually their parameters for each NEGFC subcase. Finally, we considered the subcase which led to the smallest deviations between the retrieved and injected fake companion parameters. These deviations were also used as residual noise uncertainty on the retrieved parameters.
@@
We did not retrieve the parameters for neither planet nor candidate in the F187N data, as it is not detected at a significant level in the F187N roll subtracted image. For all other cases, namely planet in the F187N and F480M datasets, and planet in the F480M dataset, we noted that the Hessian figure of merit outperformed the other two figures of merit in terms of accuracy of the inferred astrometry. It also retrieved better the contrast of the injected companions, compared to the sum figure of merit which tended to overestimate their flux, likely due to residual extended signals. For planet in the F480M data, the Hessian figure of merit is the only metric which did not significantly underestimate the radial separation of the planet compared to its expected location from the orbital fits presented in Wang et al. (2021b). The poorer performance of other metrics may be assigned to diffuse inner disk emission (see e.g. the IPCA images in Fig. 1), pushing the inferred centroid towards the center. Minimizing the determinant of the Hessian matrix is also particularly appropriate for planet c, as it involves a lesser sensitivity to the exactitude of the disk model inferred with NEGFD, that is subtracted from the cube before inferring the parameters of the planet. Using set to 2 or 3 led to consistent results for the Hessian figure of merit, while was more prone to yielding outlier values.
@@
Final uncertainties combine the residual noise uncertainties estimated with the positive fake companion injection test in quadrature with systematic uncertainties for and PA, and with photon noise uncertainties for . We considered systematic astrometric uncertainties of 6 mas and 2 mas, for the F187N and F480M data, respectively, based on the JWST user manual. The uncertainties associated to the stellar flux are negligible in comparison to the other contributions. Results are reported in the ‘Roll subtraction’ column of Table 2.
@@
As a cross check, we looked at the roll-subtracted images obtained after subtraction of both the disk model (NEGFD) and of the planet signals using the values retrieved by NEGFC. This is shown in the last column of Fig. 8. We see that the residuals at the expected (circled) location of the planets is consistent with the local noise level.
| Parameter | Roll subtraction | IPCA | Final | Prediction |
@@
| Separation | ||||
| PA | ||||
| Contrast | ||||
| Flux | ||||
@@
| Separation | – | |||
| PA | – | |||
| Contrast | – | |||
| Flux | – | |||
@@
| Separation | – | |||
| PA | – | |||
| Contrast | – | |||
| Flux | – | |||
@@
| Separation | ||||
| PA | ||||
| Contrast | ||||
| Flux | ||||
@@
| Separation | ||||
| PA | ||||
| Contrast | ||||
| Flux | ||||
Notes: Astrometric and contrast/flux predictions based on orbital fits and best-fit atmospheric models presented in Wang et al. (2021b) and M19. The contrast value reported for in M19 is corrected from the expected change in total intensity scattering efficiency based on the difference in PA with respect to the 2019 epoch. Radial separation in mas. Position angle measured east of north in . Contrast ratio with respect to the star and unresolved inner disk. Spectral flux density in Wmm.
@@ C.2 Direct NEGFC on the IPCA images
We also considered NEGFC performed directly in the IPCA images, hereafter NEGFC+IPCA, assuming self- and over-subtraction effects are (close to) fully corrected in those images. Since the IPCA images also recovered significant extended signals. It is unclear what the optimal figure of merit should be for protoplanet signals from close to, or overlapping with, complex extended signals, such as an inclined inner disk or a spiral accretion stream. Thus the major source of uncertainty for the parameters of the protoplanets for NEGFC+IPCA is rather associated to the correct assumption to be made in terms of the contribution from underlying extended signals at the location of the protoplanets. We therefore considered a range of figures of merit and associated parameters, corresponding to using either the sum, standard deviation or Hessian-matrix determinant figures of merit, for a source position corresponding to either the predicted location based on Wang et al. (2021b) and M19 orbital fits, or the visual location of a local maximum in intensity in the image. We considered two subcases corresponding to using either 1-FWHM or 2-FWHM apertures for the sum and standard deviation figures of merit, and three subcases for using either = 1, 2 or 3 for the Hessian-based figure of merit. We thus retrieved planet parameters for 14 cases in total, for each protoplanet in each dataset.
@@
Among the 14 cases, the ones considered in our estimate of final planet parameters were vetted visually upon inspection of the IPCA image after subtraction of a negative planet model with parameters found by the different figures of merit. We only considered visually pleasing results after subtraction of the estimated protoplanet parameters from the image, considering our prior knowledge of the disk (e.g., inner and outer disk geometry) such that valid estimates were typically the ones removing the local intensity peak while not creating a significant hole within the surrounding extended signal distribution. The selection of different cases depended on the considered protoplanet and filter, as each image and protoplanet location are affected by more or less surrounding circumstellar signals. We then considered the median of the results obtained for the cases validated visually, and adopted the standard deviation of these results as the uncertainty associated to the contamination by surrounding extended signals. As for the classical NEGFC approach, we added in quadrature systematic astrometric uncertainties and photon noise uncertainties to the uncertainties associated to contamination by surrounding extended signals. Our results are reported in the ‘IPCA’ column of Table 2.
@@
Given the overlap between the signals of planet , candidate , the spiral accretion stream and inner disk signals, we adopted an iterative joint-fitting strategy to estimate the parameters for planet and candidate d. We started by estimating the parameters from , as it is more easily dissociated from surrounding extended signals, estimated the parameters for d in the image without the signal of (similar to Fig. 1c), then re-estimated the parameters for this time in an image where the estimated contribution from d was removed, and so on. This procedure converged to stable estimated parameters for both sources within five iterations. We conservatively adopted for both sources the astrometric uncertainties corresponding to the largest uncertainty found among the two sources.
@@
As a safety check, we inspected the IPCA images after subtraction of the optimal parameters found for the protoplanets with the direct NEGFC procedure, using the image cube where the optimal outer disk model was subtracted (NEGFD), to ensure that the estimates did not over- nor underestimated the flux.

### v2

@@ A.3 Reliability of the IPCA
We tested the IPCA method on a fiducial dataset in order to illustrate its effectiveness at producing unbiased images of extended circumstellar signals, and to validate all our conclusions regarding the flux and morphology of circumstellar signals recovered in our IPCA images of PDS 70 in this way. For this test, we considered the only other JWST/NIRCam dataset obtained in the F187N filter with the NRCB1 detector and SUB64P subarray that was publicly available at the time of this study: an observation of reference star HD 135067 that was obtained on 2 February 2023 (Program 1902). This source does not have any known resolved circumstellar emission. We directly downloaded the calibrated i2d images from the MAST archive. While the number of integrations and the number of groups per integration were similar to the PDS 70 observation, the main difference was the adoption of a three-point dither pattern strategy instead of a five-point dither pattern for the observation of PDS 70. This results in a higher sensitivity to spatial undersampling effects, which are expected to be particularly prominent near the core of the PSF. As a consequence, the results of the test presented in this section should be considered as a conservative lower limit on the actual expected performance of IPCA on the PDS 70 dataset.
@@
We designed a toy model for the circumstellar signals of PDS 70 that was composed of an outer disk, an inner disk, a spiral-like signal, and three point-source injections corresponding to protoplanets and , and candidate . For the outer disk, we considered the optimal model described in Appendix B. For the inner disk, we considered similar parameters as presented in Keppler et al. (2018), with the peak of the emission concentrated within 5 au. The two disk components are shown in Fig. 6a. For the spiral-like signal, we considered a similar trace as observed in our F187N images of PDS 70 after subtraction of the optimal outer-disk model (Fig. 6b). For the protoplanets, we considered similar positions and contrasts as inferred with NEGFC in Appendix C (Fig. 6c). The sum of all components of the model is shown in Fig. 6d. To create and inject the fiducial model into the reference cube, we relied on routines from the fm (forward-modeling) subpackage of VIP, in particular, its implementation of the Grenoble RAdiative TransfER tool (Augereau et al. 1999; Milli et al. 2019) for the inner-disk model, its trace and fake-companion injection routines. Before injection in the reference cube data, the model was scaled to a similar contrast with respect to the reference star as the circumstellar signals with respect to PDS 70. The signals were injected at two different angles separated by 5.0 deg in the two sets of three dithered images (i.e., the same roll-angle difference as in the PDS 70 observation).
@@
Finally, we ran IPCA on this fiducial dataset using the same reduction parameters as were used to produce the F187N images of PDS 70. The algorithm converged within 400 iterations, resulting in the image shown in Fig. 6e. To facilitate the comparison with the ground-truth injected model, we highlight the location of the injected protoplanets with dashed circles. Stellar residual artifacts can likely be assigned to small differences in the core of the PSF and spatial undersampling effects, which are more prominent when using a three-point instead of a five-point dither pattern. Apart from theses residuals, we note a satisfactory recovery of both the morphology and flux level of the injected circumstellar signals. This makes us confident of the results obtained with IPCA on the PDS 70 dataset reported in this paper.
@@ Appendix B Optimal disk model found with NEGFD
To be able to extract unbiased photometry for the protoplanets, the expected contribution of the disk must be removed. Our goal in this work is not a full modeling of the disk, as this would involve a combined SED and disk-image fit (e.g., Keppler et al. 2018; Portilla-Revelo et al. 2022, 2023). Therefore, we decided to rely on the latest radiative transfer models of the disk produced with MCMax3D (Min et al. 2009) and presented in Portilla-Revelo et al. (2022, 2023), as they match these combined constraints, and only allow these models to vary to a small extent so that they are still compatible with ALMA and SPHERE polarized-intensity constraints. Namely, we allowed (i) the minimum grain size in the grain size distribution to vary between 0.001 and 0.05 m; (ii) the settling parameter to vary between 0.001 m and 0.01 m; (iii) spatial and flux scaling of the model image to vary within a small range around unity; and (iv) small linear (subpixel) and azimuthal (subdegree) shifts with respect to the center of the image, as these parameters are essentially constrained by imaging. Small variations in these parameters can lead to noticeable changes in the model images and therefore, to significant residuals after subtraction of the model. We produced a grid of ten disk models in (two explored values) and (five explored values), and searched for the optimal model with a Nelder-Mead algorithm by interpolating model images in log-space and including free parameters for scaling and shifts. The model images were produced at the same pixel scale as the F187N and F480M images, and they were smoothed by convolution with the observed PSF. The optimal model was then found by minimizing the absolute residuals after subtraction of the model from the observed images. This procedure, which we refer to as the negative fake disk technique (NEGFD), is implemented in VIP as of version 1.6.0.
@@
The optimal disk model was found by minimizing the sum of absolute intensity residuals in a binary mask that encompassed the location of the outer disk while excluding the location of both planets (2-FWHM aperture exclusion masks). We considered two potential masks. The first mask included the whole forward-scattering (i.e., brighter) side of the disk, and the second map only included the southwest part of the disk, anticipating excess signals toward the west (near planet c) and northwest part of the disk (spiral-like feature) based on previous images of the disk (Wang et al. 2020; Juillard et al. 2022). As our tests using the first mask led to a mix of strong positive and negative residuals, we only consider the results obtained by minimizing residuals in the second mask, shown in Fig. 7c, in the rest of this work. In practice, we identified the optimal radiative transfer disk model using the F480M data because of the higher signal-to-noise ratio of the disk in these data. The optimal and values were found to be 0.001 m and 0.01, respectively, meaning that the optimal model is in between the radiative transfer models considered in Portilla-Revelo et al. (2022) and Portilla-Revelo et al. (2023). These parameters were then used to make a disk model prediction at 1.87 m. The best F187N and F480M models are shown in Fig. 7a and b, respectively.
@@
Figs. 8b and e show the roll subtraction images obtained after subtracting the optimal disk model found by NEGFD from the original images (i.e., before PSF subtraction) for the F187N and F480M datasets, respectively. These are compared to the original roll subtraction images (Figs. 8a and d) using the same intensity scale. The disk subtraction performs well, in particular, for the southwest part of the disk, where the residuals reach a similar level as the residual noise level in the image. Subtraction of the optimal disk model clearly highlights the presence in the F480M image of the arm-like feature characterized in Juillard et al. (2022) in VLT/SPHERE images. This feature is indicated with a thick arrow. We also checked how a similar NEGFD procedure performed on the IPCA images obtained with both filters by fixing the and to the optimal values found by NEGFC+roll subtraction. The results are shown in Fig. 1b and e, and they are discussed in Sec. 3.1.
@@ Appendix C NEGFC retrieval of planet parameters
The NEGFC technique consists of finding the optimal parameters (radial separation, position angle, and contrast with respect to the star) of a directly imaged companion through the injection of fake companions with a negative flux in the input image cube (i.e., before post-processing), such that residuals are minimized in the post-processed image around the location of the planet. This forward-modeling approach is necessary to avoid the parameter estimates for the companion to be affected by self- and oversubtraction effects inherent to the post-processing algorithm used.
We used the Nelder-Mead minimization algorithm implemented in VIP and described in Wertz et al. (2017) to estimate the parameters for the protoplanets. This NEGFC implementation offers different options in terms of figure of merit to be used in the processed image after injection of negative fake companions to identify the optimal radial separation , position angle PAp, and contrast of a given planet, namely minimizing (i) the sum of absolute intensity residuals in an aperture encompassing the companion, (ii) the standard deviation of intensity residuals in an aperture encompassing the companion, or (iii) the sum of absolute determinant values of the Hessian matrix calculated for each of the pixels surrounding the planet location. The latter option is equivalent to minimizing the local absolute curvature, and it is more appropriate for the extraction of point-source astrometry and photometry in the presence of underlying extended signals (e.g., Quanz et al. 2015). We implemented it for this work and made it available in VIP as of version 1.6.0. We detail two different approaches below that involve NEGFC that we tested and compared to retrieve optimal astrometry and photometry for the protoplanets.
@@ C.1 Classical forward-modeling NEGFC
We first considered the classical forward-modeling NEGFC approach as described in Lagrange et al. (2010) or Wertz et al. (2017), combined with roll subtraction, hereafter referred to as NEGFC+roll. We applied NEGFC+roll to the image cube where the optimal disk model found by NEGFD was removed (Appendix B). Regarding the choice of figure of merit, we note that different figures of merit lead to more or less reliably retrieved planet parameters depending on the S/N of the companion, the nature of the local noise, and contamination by residual extended signals (e.g., inner disk or spiral accretion stream signal). We therefore determined the optimal figure of merit on a case-by-case basis for each planet in each of our two datasets as follows. We used NEGFC to derive first estimates on the parameters of the companion with each of the three figures of merit, considering two subcases for the sum and standard deviation, corresponding to their calculation in either 1-FWHM or 2-FWHM size apertures, and three subcases for the Hessian-matrix determinant figure of merit, corresponding to its calculation with set to 1, 2, or 3. Then we injected 360 fake (positive) companions with the radial separation and contrast inferred by NEGFC, 1 apart from each other in separate copies of the original datasets, and individually retrieved their parameters for each NEGFC subcase. Finally, we considered the subcase that led to the smallest deviations between the retrieved and injected fake companion parameters. These deviations were also used as residual noise uncertainty on the retrieved parameters.
@@
We did not retrieve the parameters for planet or candidate in the F187N data because they are not detected at a significant level in the F187N roll-subtracted image. For all other cases, namely planet in the F187N and F480M datasets and planet in the F480M dataset, we note that the Hessian figure of merit outperformed the other two figures of merit in terms of accuracy of the inferred astrometry. It also retrieved better the contrast of the injected companions than the sum figure of merit, which tended to overestimate their flux, likely due to residual extended signals. For planet in the F480M data, the Hessian figure of merit is the only metric that did not significantly underestimate the radial separation of the planet compared to its expected location from the orbital fits presented in Wang et al. (2021b). The poorer performance of other metrics may be assigned to diffuse inner-disk emission (see, e.g., the IPCA images in Fig. 1), pushing the inferred centroid toward the center. Minimizing the determinant of the Hessian matrix is also particularly appropriate for planet c because it involves a lower sensitivity to the exactitude of the disk model inferred with NEGFD, which is subtracted from the cube before inferring the parameters of the planet. set to 2 or 3 led to consistent results for the Hessian figure of merit, while was more prone to yielding outlier values.
@@
The final uncertainties combine the residual noise uncertainties estimated with the positive fake companion injection test in quadrature with systematic uncertainties for and PAp and with photon-noise uncertainties for . We considered systematic astrometric uncertainties of 6 mas and 2 mas, for the F187N and F480M data, respectively, based on the JWST user manual. The uncertainties associated with the stellar flux are negligible in comparison to the other contributions. The results are reported in the ‘Roll subtraction’ column of Table 2.
@@
As a cross check, we examined the roll-subtracted images obtained after subtraction of both the disk model (NEGFD) and of the planet signals using the values retrieved by NEGFC. This is shown in the last column of Fig. 8. The residuals at the expected (circled) location of the planets are consistent with the local noise level.
| Parameter | Roll subtraction | IPCA | Final | Prediction(a) |
@@
| Separation(b) | ||||
| PA(c) | ||||
| Contrast(d) | ||||
| Flux(e) | ||||
@@
| Separation(b) | – | |||
| PA(c) | – | |||
| Contrast(d) | – | |||
| Flux(e) | – | |||
@@
| Separation(b) | – | |||
| PA(c) | – | |||
| Contrast(d) | – | |||
| Flux(e) | – | |||
@@
| Separation(b) | ||||
| PA(c) | ||||
| Contrast(d) | ||||
| Flux(e) | ||||
@@
| Separation(b) | ||||
| PA(b) | ||||
| Contrast(c) | ||||
| Flux(e) | ||||
Notes: (a)Astrometric and contrast or flux predictions based on orbital fits and best-fit atmospheric models presented in Wang et al. (2021b) and M19. The contrast value reported for in M19 is corrected for the expected change in total intensity-scattering efficiency based on the difference in PA with respect to the 2019 epoch. (b)Radial separation in mas. (c)Position angle measured east of north in . (d)Contrast ratio with respect to the star and unresolved inner disk. (e)Spectral flux density in Wmm-1.
@@ C.2 Direct NEGFC on the IPCA images
We also considered NEGFC performed directly on the IPCA images, hereafter NEGFC+IPCA, assuming that self- and oversubtraction effects are (close to) fully corrected for in these images because the IPCA images also recovered significant extended signals. It is unclear what the optimal figure of merit should be for protoplanet signals from close to or overlapping with complex extended signals, such as an inclined inner disk or a spiral accretion stream. The main source of uncertainty for the parameters of the protoplanets for NEGFC+IPCA is therefore rather associated with the correct assumption to be made in terms of the contribution from underlying extended signals at the location of the protoplanets. We therefore considered a range of figures of merit and associated parameters, corresponding to using either the sum, standard deviation, or Hessian-matrix determinant figures of merit for a source position corresponding to either the predicted location based on Wang et al. (2021b) and M19 orbital fits or the visual location of a local maximum in intensity in the image. We considered two subcases corresponding to using either 1-FWHM or 2-FWHM apertures for the sum and standard deviation figures of merit, and three subcases for using either = 1, 2, or 3 for the Hessian-based figure of merit. We thus retrieved planet parameters for 14 cases in total for each protoplanet in each dataset.
@@
These 14 cases we considered in our estimate of the final planet parameters were vetted visually upon inspection of the IPCA image after subtracting a negative planet model with parameters determined by the different figures of merit. We only considered visually pleasing results after subtracting the estimated protoplanet parameters from the image considering our prior knowledge of the disk (e.g., inner and outer-disk geometry), such that valid estimates typically were those from which the local intensity peak was removed while not creating a significant hole within the surrounding extended signal distribution. The selection of different cases depended on the considered protoplanet and filter because each image and protoplanet location are affected by more or fewer surrounding circumstellar signals. We then considered the median of the results obtained for the cases validated visually and adopted the standard deviation of these results as the uncertainty associated with the contamination by surrounding extended signals. For the classical NEGFC approach, we added systematic astrometric uncertainties and photon noise uncertainties in quadrature to the uncertainties associated with contamination by surrounding extended signals. Our results are reported in the ‘IPCA’ column of Table 2.
@@
Because of the overlap between the signals of planet , candidate , the spiral accretion stream, and the inner disk signals, we adopted an iterative joint-fitting strategy to estimate the parameters for planet and candidate d. We started by estimating the parameters from because it is more easily dissociated from surrounding extended signals, estimated the parameters for d in the image without the signal of (similar to Fig. 1c), then reestimated the parameters for , this time, in an image from which the estimated contribution of d was removed, and so on. This procedure converged to stable estimated parameters for both sources within five iterations. We conservatively adopted for both sources the astrometric uncertainties corresponding to the largest uncertainty found among the two sources.
@@
As a safety check, we inspected the IPCA images after subtracting the optimal parameters found for the protoplanets with the direct NEGFC procedure using the image cube from which the optimal outer-disk model was subtracted (NEGFD) to ensure that the estimates did not over- or underestimated the flux.